A steel ball with a mass of 3 kg is applied to a large thick wall at a speed of 10 m / s
crashed at an angle of 60 ° to the plane. The ball bounces at the same speed and at the same angle. The ball
The ball was served from the wall while the wall was touched for 0.2 seconds
Find the average power
1
Expert's answer
2020-05-04T12:47:00-0400
"F=\frac{2mv\sin{60}}{\Delta t}"

"P=Fv\sin{60}=\frac{2mv^2\sin^2{60}}{\Delta t}"

"P=\frac{2(3)(10)^2\sin^2{60}}{0.2}=2250\ W"